# Fairness-Regulated Dense Subgraph Discovery

### v1

@@ 1 Introduction
Ensuring that data-driven algorithms do not disproportionately impact different population subgroups is a key challenge in human-centered applications of artificial intelligence. The field of algorithmic fairness [1, 2, 3, 4] focuses on designing fairness-enhancing mechanisms to mitigate algorithmic bias. Fairness considerations have been formulated as optimization problems in both supervised and unsupervised learning (see the surveys [5, 6] and references therein). In contrast, it remains a nascent area of research in graph mining, where fairness considerations are also important [7]. To date, only few graph mining problems have been analyzed under the lens of algorithmic fairness.
@@
Dense subgraph discovery (DSD) is one key graph mining primitive that aims to extract subgraphs with high internal connectivity from a given graph (see the survey [8] and references therein). DSD finds widespread applications ranging from mining social media trends [9], detecting patterns in gene annotation graphs [10], and spotting fraud in e-commerce and financial networks [11, 12, 13]. A popular formulation used for extracting dense subgraphs maximizes the average induced degree [14]. This is because the problem can be solved exactly using maximum-flows, or approximated efficiently at scale using greedy vertex-peeling algorithms [15, 16, 17], or convex optimization algorithms [18, 19, 20], with guarantees on the sub-optimality of the solution.
@@
In this paper, we study the problem of extracting the densest subgraph from a given graph that meets a target fairness criterion. Such considerations can arise in DSD when the vertices of the graph (e.g., corresponding to a web or social network) are annotated with sensitive information pertaining to an individual’s gender/race/religion/political leaning, etc., which partitions the vertex set into different population subgroups. The fair dense subgraph problem then corresponds to locating a dense subgraph with adequate levels of representation of each subgroup in the extracted subset. Simply applying a pre-existing algorithm for DSD can fail in this regard. Indeed, studies on real-world graphs [21, 22] have revealed that existing methods return subgraphs which typically exhibit strong homophily, with little to no diversity in the composition of the vertex attributes. Motivating examples of fairness in DSD include selection of diverse teams from a social network of professional contacts [23, 24], and recommending balanced content to social media users that spans the various views of the individuals comprising the network, thus mitigating polarization [25].
@@
While DSD is a well-studied topic, extracting fair dense subgraphs has received limited attention. Although the recent works of [21, 22, 26] have initiated progress in this area, they have some inherent limitations. Notably, their formulations are NP–hard in the worst-case, necessitating approximation algorithms. These methods also struggle with rigid or hard-to-set fairness criteria. Another key issue is the inherent trade-off between subgraph density and fairness, which is difficult to analyze and has not been formally explored due to the NP-hardness of existing approaches (see 2 section for details).
@@
Contributions: In this paper, we focus on the setting where the vertices of the graph are divided in two subgroups, protected and unprotected. Our objective is to overcome the limitations of prior work by (i) proposing new, tractable (polynomial-time solvable) formulations which are (ii) capable of generating a spectrum of fair dense subgraphs with varying levels of fairness. This approach enables users to analyze the trade-off between subgraph density and fair representation of the protected group, and quantify the price of fairness, i.e., the density loss required to meet a target fairness level. We point out that these features are absent in the formulations of [21, 22, 26]. Our key technical idea is to promote fairness in DSD through a pair of judiciously chosen regularizers that create tractable problems and establish regularization paths representing various fairness levels. To summarize:
@@
  • We analyze the inherent trade-off between subgraph density and target fairness for a difficult example using the price of fairness metric. Our results indicate that enhancing fairness can significantly reduce density, regardless of the algorithm used.
@@ 2 Related Work
Densest subgraph and it variants: Given a graph $\mathcal{G}$, the classic Densest Subgraph (DSG) problem [14] aims to extract a dense vertex subset $\mathcal{S}\subseteq\mathcal{V}$ that maximizes the average (induced) degree. DSG can be solved exactly in polynomial time via maximum flows, but this is computationally demanding for large graphs. In practice, a scalable greedy algorithm that iteratively peels vertices is used, providing a $1/2$-factor approximation [27, 15]. The algorithm was generalized in the recent work of [16], which proposed a multi-stage greedy algorithm Greedy++ that exhibits fast convergence to a near-optimal solution of DSG without relying on flows. The follow-up work of [17] provided a performance analysis of Greedy++, and demonstrated that it can attain $(1-\epsilon)$-approximate solutions of DSG in $O(1/\epsilon^{2})$ iterations. Building on this approach, a general peeling-based framework Super-Greedy++ was introduced in [17] for obtaining $(1-\epsilon)$-approximate solutions for the broader problem class of computing densest supermodular subsets, which can be expressed as $\max_{\mathcal{S}\subseteq\mathcal{V}}\{F(\mathcal{S})/|\mathcal{S}|\}$, where $F(\mathcal{S})$ is a monotone, normalized, supermodular function. Note that DSG is a special case, with $F(\mathcal{S})=e(\mathcal{S})$. More generally, it was shown in [18] that Super-Greedy++ exhibits convergence to an optimal dense decomposition vector of $f$, which corresponds to a partition of of $\mathcal{V}$ into a finite collection of subsets $\mathcal{S}_{1},\cdots,\mathcal{S}_{k}$, such that each for each $i\in\{1,\cdots,k\}$, $\mathcal{S}_{i}$ is the unique maximal subset that maximizes $\lambda_{i}=(F(\mathcal{S}_{i}\cup\mathcal{X}_{i-1})-F(\mathcal{X}_{i-1}))/|\mathcal{S}_{i}|$ over $\mathcal{V}\setminus\mathcal{X}_{i-1}$, where $\mathcal{X}_{i-1}:=\cup_{j<i}S_{j}$. The value of $\lambda_{1}$ corresponds to density of the densest supermodular subset, whereas subsequent values monotonically decrease, i.e., $\lambda_{1}>\cdots>\lambda_{k}$. For DSG, we refer to such a decomposition as a dense decomposition on a graph $\mathcal{G}$.
@@
A separate line of research [28, 20, 19] for developing fast approximation algorithms for DSG has blossomed around applying convex-optimization algorithms for solving a tight linear programming (LP) relaxation of DSG originally proposed in [15]. The work of [28] considered a quadratic programming problem derived from the dual of the LP relaxation of [15], and applied the Frank-Wolfe algorithm [29] for solving it. Recently, the authors of [20] developed faster algorithms for DSG and dense decompositions on graphs by solving the dual LP relaxation via accelerated proximal gradient [30]. This was later improved upon by [19] using tools from area convexity [31] and accelerated random coordinate descent [32] to develop faster approximation algorithms for DSG and dense decompositions on graphs, respectively.
@@
An important variant of DSG is finding the $k$-core decomposition of a graph [33]. A $k$-core is the maximal subgraph (by size) where every vertex has an induced degree of at least $k$. The largest value of $k$ for which such a subgraph exists is known as the degeneracy, and the corresponding subgraph is known as the max-core. Extracting the max-core is tantamount to finding the vertex subset which maximizes the minimum induced degree. A slight modification of the greedy peeling algorithm for DSG exactly solves this problem.
@@
Recently, it has been shown [34] that the solution of DSG and the $k$-core both correspond to special cases of a more general framework known as the Generalized Mean Densest Subgraph (GMDSG) problem. This framework is parameterized by a single parameter $p$, based on computing generalized means of degree sequences of a subgraph, with DSG and $k$-core corresponding to the choices of $p=1$ and $p=-\infty$, respectively. Hence, by varying $p$, one can extract a family of dense subgraphs which obey different notions of density. For $p\geq 1$, GMDSG can be solved optimally in polynomial-time via maximum flows, or approximated via a generalized greedy peeling algorithm. Additionally, it was shown in [35] that for $p<1$, GMDSG is NP–hard for $p$ in the range $p\in(-1/8,0)\cup(0,1/4)$. The authors of [35] also showed that the solutions of DSG and max-core provide a $1/2$-approximation for GMDSG when $p<1$.
@@
Fairness in DSD: Subgraphs extracted via DSG or its variants are not guaranteed to fulfill a pre-specified fairness criterion (regardless of the algorithmic approach employed), as the respective formulations do not explicitly take into account any protected attributes that the vertices may possess. This issue was recently brought up in [21, 22, 26], which considered the task of extracting dense subgraphs that satisfy a group-fairness criterion. In [21, 22], each vertex is a member of either a protected or an unprotected group, and the goal is to extract the densest subgraph with an equal number of vertices from both groups (perfect balance). The follow-up work of [26] proposed a pair of fairness promoting formulations which can handle multiple vertex groups. The first [26, Problem 1] corresponds to an extension of [21, 22] that aims to find the densest subgraph subject to the constraint that the number of vertices from each group in the solution does not exceed a fixed fraction of the solution size. The second formulation [26, Problem 2] concerns finding the densest subgraph while ensuring that at least a pre-specified number of vertices from each group are present in the solution. A separate work [42] considered a variation of the fair DSG problem where the edge relationships (as opposed to vertices) in a graph are modeled as belonging to different groups. The authors of [42] proposed formulations for extracting fair dense subgraphs containing exactly, at most, and at least a certain number of edges from each edge type.
@@ 3 Problem Statement
Given a graph $\mathcal{G}$ with designated protected and unprotected subgroups of vertices, let ${\mathbf{e}}_{p}\in\{0,1\}^{n}$ denote the binary indicator vector of the protected subset $\mathcal{S}_{p}$. Define the regularizer $r_{1}({\mathbf{x}}):=({\mathbf{e}}_{p}^{\top}{\mathbf{x}})/{\mathbf{e}}^{\top}{\mathbf{x}}$ and consider the following problem
${\mathbf{x}}^{*}(\lambda)\in\operatorname*{argmax}_{{\mathbf{x}}\in\{0,1\}^{n}}~{}\biggl{\{}g({\mathbf{x}},\lambda):=\rho({\mathbf{x}})+\lambda\cdot r_{1}({\mathbf{x}})\biggr{\}}.$ (1)
We designate the above formulation as the first Fairness-Aware Densest Subgraph (FADSG-I) problem. Here, $\lambda\geq 0$ is the regularization parameter. The objective function $g({\mathbf{x}},\lambda)$ consists of two terms - the first term $\rho({\mathbf{x}})$ corresponds to subgraph density, whereas for $\lambda>0$, $r_{1}({\mathbf{x}})$ promotes the inclusion of vertices from the protected subset $\mathcal{S}_{p}$. This can be seen by equivalently expressing $r_{1}(\cdot)$ as the set function
@@
It is evident that $0\leq r_{1}(\mathcal{S})\leq 1$, with the upper bound being attained for any subset $\mathcal{S}\subseteq\mathcal{S}_{p}$. Meanwhile, $r_{1}(\mathcal{S})=0$ iff $\mathcal{S}$ does not contain any protected vertices. Thus, larger values of $r_{1}(\cdot)$ correspond to vertex subsets containing a greater fraction of protected vertices. Note that $r_{1}(\mathcal{S})=1/2$ corresponds to the case of perfect balance [21, 22], with equal number of protected and unprotected vertices in $\mathcal{S}$. By varying $\lambda$, we allow $r_{1}(\cdot)$ to span a range of values, which affords flexibility in extracting dense subgraphs exhibiting varying fairness levels.
@@
When $\lambda=0$, FADSG-I reduces to DSG, the solution of which is not guaranteed to contain any protected vertices. As $\lambda$ is increased, greater emphasis is placed on maximizing $r_{1}({\mathbf{x}})$, enhancing the inclusion of protected vertices. For $\lambda\rightarrow\infty$, $r_{1}({\mathbf{x}}^{*}(\lambda))\rightarrow 1$, and the solution of problem (1) corresponds to the densest subset of $\mathcal{S}_{p}$. Clearly, by varying $\lambda$ in the interval $[0,\infty)$, the solution of (1) spans a spectrum which lies between these two extremes. By appropriately selecting $\lambda$, desired representation levels of protected subset can be achieved. Hence, in contrast to [21, 22, 26], our group-fairness formulation (1) affords much greater flexibility in trading-off subgraph density in exchange for protected subset representation.
@@
It is worth pointing out that in certain scenarios, the solution of FADSG-I may not provide satisfactory levels of representation of the protected subset. Although the regularizer $r_{1}(\cdot)$ promotes the inclusion of protected vertices in the solution of (1), it does not explicitly take into account the fraction of such vertices that are represented. Even for (infinitely) large values of $\lambda$ in (1), the maximal level of representation attainable corresponds to the densest subset of $\mathcal{S}_{p}$. In general, this does not imply that a fixed fraction of the vertices of $\mathcal{S}_{p}$ will be included in the solution. These considerations motivate the following alternative
${\mathbf{x}}^{*}(\lambda)\in\operatorname*{argmax}_{{\mathbf{x}}\in\{0,1\}^{n}}~{}\biggl{\{}h({\mathbf{x}},\lambda):=\rho({\mathbf{x}})-\lambda\cdot r_{2}({\mathbf{x}})\biggr{\}},$ (3)
@@
where $r_{2}({\mathbf{x}}):=(\lVert{\mathbf{x}}-{\mathbf{e}}_{p}\rVert_{2}^{2})/{\mathbf{e}}^{\top}{\mathbf{x}}$ is a regularizer and $\lambda\geq 0$ is a regularization parameter. The above problem is designated as FADSG-II, and differs from FADSG-I in the choice of the regularizer. Note that the numerator of $r_{2}(\cdot)$ can be viewed as the Hamming distance between the sets $\mathcal{S}$ and $\mathcal{S}_{p}$, since $\lVert{\mathbf{x}}-{\mathbf{e}}_{p}\rVert_{2}^{2}=\sum_{i\in\mathcal{V}}({\mathbf{x}}(i)-{\mathbf{e}}_{p}(i))^{2}$ corresponds to the number of disagreements between $\mathcal{S}$ (with indicator vector ${\mathbf{x}}$) and $\mathcal{S}_{p}$ (with indicator vector ${\mathbf{e}}_{p}$). Hence, for a given value of $\lambda>0$, the objective function of (3) aims to maximize subgraph density while penalizing the distance between the extracted and the protected subgraphs. As $\lambda\rightarrow\infty$, $r_{2}({\mathbf{x}}^{*}(\lambda))\rightarrow 0$, and the solution ${\mathbf{x}}^{*}(\lambda)$ coincides with the entire protected subset itself. Building on this observation, consider the case where the target fairness requirement stipulates that $50\%$ of the vertices in the protected subset be included in $\mathcal{S}$. We have the following result (see Appendix A.1 for proof).
@@
More generally, although problems (1) and (3) promote different notions of fairness, it is interesting to note that the two formulations are not completely unrelated. In particular, we have shown that the following relationship holds (see Appendix A.2 B for the proof).
@@
$1-r_{1}(\mathcal{S})\leq r_{2}(\mathcal{S})$.
The implication of the above result is that whenever $r_{2}(\mathcal{S})\in[0,1]$, we obtain a non-trivial lower-bound on $r_{1}(\cdot)$. Since providing explicit representation of the protected subset is absent in the fairness formulations of [21, 22, 26], FADSG-II goes beyond what is attainable using the prior art.
Tractability: A limitation of the formulations of [21, 22, 26] is that they are NP–hard in the worst-case. Since optimality cannot be guaranteed, this implies that the quality of the solution obtained by their approximation algorithms may not be satisfactory in terms of the loss in density incurred for a target fairness level. In contrast, it turns out that problems (1) and (3) can always be solved exactly in polynomial-time, irrespective of $\lambda$ (see Appendix B for details). This implies that by varying $\lambda$, we can trace the Pareto-optimal boundary between the two components comprising the objective function (i.e., density and fairness) of both problems (1) and (3). This enables us to reveal the underlying trade-off between density and fairness on real-world datasets. As mentioned before, the intrinsic hardness of the formulations in [21, 22, 26] does not allow for such a feature.

### v2

# Fairness-Aware Dense Subgraph Discovery

@@ 1 Introduction
Ensuring that data-driven algorithms do not disproportionately impact different population subgroups is a key challenge in human-centered applications of artificial intelligence. The field of algorithmic fairness (Dwork et al., 2012; Feldman et al., 2015; Kleinberg et al., 2018; Holstein et al., 2019) focuses on designing fairness-enhancing mechanisms to mitigate algorithmic bias. Fairness considerations have been formulated as optimization problems in both supervised and unsupervised learning (see the surveys Friedler et al. (2019); Chouldechova & Roth (2020) and references therein). In contrast, it remains a nascent area of research in graph mining, where fairness considerations are also important (Dong et al., 2023). To date, only few graph mining problems have been analyzed under the lens of algorithmic fairness.
@@
Dense subgraph discovery (DSD) is one key graph mining primitive that aims to extract subgraphs with high internal connectivity from a given graph (see the survey Lanciano et al. (2024) and references therein). DSD finds widespread applications ranging from mining social media trends (Angel et al., 2012), detecting patterns in gene annotation graphs (Khuller & Saha, 2009), and spotting fraud in e-commerce and financial networks (Hooi et al., 2016; Li et al., 2020; Chen & Tsourakakis, 2022). A popular formulation used for extracting dense subgraphs maximizes the average induced degree (Goldberg, 1984). This is because the problem can be solved exactly using maximum-flows, or approximated efficiently at scale using greedy vertex-peeling algorithms (Charikar, 2000; Boob et al., 2020; Chekuri et al., 2022), or convex optimization algorithms (Harb et al., 2023; Nguyen & Ene, 2024; Harb et al., 2022), with guarantees on the sub-optimality of the solution.
@@
In this paper, we study the problem of extracting the densest subgraph from a given graph that meets a target fairness criterion. Such considerations can arise in DSD when the vertices of the graph (e.g., corresponding to a web or social network) are annotated with sensitive information pertaining to an individual’s gender/race/religion/political leaning, etc., which partitions the vertex set into different population subgroups. The fair dense subgraph problem then corresponds to locating a dense subgraph with adequate levels of representation of each subgroup in the extracted subset. Simply applying a pre-existing algorithm for DSD can fail in this regard. Indeed, studies on real-world graphs (Anagnostopoulos et al., 2020; 2024) have revealed that existing methods return subgraphs which typically exhibit strong homophily, with little to no diversity in the composition of the vertex attributes. Motivating examples of fairness in DSD include selection of diverse teams from a social network of professional contacts (Marcolino et al., 2013; Rangapuram et al., 2013), and recommending balanced content to social media users that spans the various views of the individuals comprising the network, thus mitigating polarization (Musco et al., 2018).
@@
While DSD is a well-studied topic, extracting fair dense subgraphs has received limited attention. Although the recent works of Anagnostopoulos et al. (2020; 2024); Miyauchi et al. (2023) have initiated progress in this area, they have some inherent limitations. Notably, their formulations are NP–hard in the worst-case, necessitating approximation algorithms. These methods also struggle with rigid or hard-to-set fairness criteria. Another key issue is the inherent trade-off between subgraph density and fairness, which is difficult to analyze and has not been formally explored due to the NP-hardness of existing approaches (see Section 2 for details).
@@
Contributions: In this paper, we focus on the setting where the vertices of the graph are divided in two subgroups, protected and unprotected. Our objective is to overcome the limitations of prior work by (i) proposing new, tractable (polynomial-time solvable) formulations which are (ii) capable of generating a spectrum of fair dense subgraphs with varying levels of fairness. This approach enables users to analyze the trade-off between subgraph density and fair representation of the protected group, and quantify the price of fairness, i.e., the density loss required to meet a target fairness level. We point out that these features are absent in the formulations of Anagnostopoulos et al. (2020; 2024); Miyauchi et al. (2023). Our key technical idea is to promote fairness in DSD through a pair of judiciously chosen regularizers that create tractable problems and establish regularization paths representing various fairness levels.
To summarize:
@@
  • We analyze the inherent trade-off between subgraph density and target fairness for a difficult example using the price of fairness measure. Our results indicate that enhancing fairness can significantly reduce density, regardless of the algorithm used.
@@ 2 Related Work
Densest subgraph and it variants: Given a graph $\mathcal{G}$, the classic Densest Subgraph (DSG) problem (Goldberg, 1984) aims to extract a dense vertex subset $\mathcal{S}\subseteq\mathcal{V}$ that maximizes the average (induced) degree. DSG can be solved exactly in polynomial time via maximum flows, but this is computationally demanding for large graphs. In practice, a scalable greedy algorithm that iteratively peels vertices is used, providing a $1/2$-factor approximation (Asahiro et al., 2000; Charikar, 2000). The algorithm was generalized in the recent work of Boob et al. (2020), which proposed a multi-stage greedy algorithm Greedy++ that exhibits fast convergence to a near-optimal solution of DSG without relying on flows. The follow-up work of Chekuri et al. (2022) provided a performance analysis of Greedy++, and demonstrated that it can attain $(1-\epsilon)$-approximate solutions of DSG in $O(1/\epsilon^{2})$ iterations. Building on this approach, a general peeling-based framework Super-Greedy++ was introduced by Chekuri et al. (2022) for obtaining $(1-\epsilon)$-approximate solutions for the broader problem class of computing densest supermodular subsets, which can be expressed as $\max_{\mathcal{S}\subseteq\mathcal{V}}\{F(\mathcal{S})/|\mathcal{S}|\}$, where $F(\mathcal{S})$ is a monotone, normalized, supermodular function. Note that DSG is a special case, with $F(\mathcal{S})=e(\mathcal{S})$. More generally, it was shown by Harb et al. (2023) that Super-Greedy++ exhibits convergence to an optimal dense decomposition vector of $f$, which corresponds to a partition of of $\mathcal{V}$ into a finite collection of subsets $\mathcal{S}_{1},\ldots,\mathcal{S}_{k}$, such that each for each $i\in\{1,\ldots,k\}$, $\mathcal{S}_{i}$ is the unique maximal subset that maximizes $\lambda_{i}=(F(\mathcal{S}_{i}\cup\mathcal{X}_{i-1})-F(\mathcal{X}_{i-1}))/|\mathcal{S}_{i}|$ over $\mathcal{V}\setminus\mathcal{X}_{i-1}$, where $\mathcal{X}_{i-1}:=\cup_{j<i}S_{j}$. The value of $\lambda_{1}$ corresponds to density of the densest supermodular subset, whereas subsequent values monotonically decrease, $\lambda_{1}>\cdots>\lambda_{k}$. For DSG, we refer to such a decomposition as a dense decomposition on a graph $\mathcal{G}$.
@@
A separate line of research (Danisch et al., 2017; Harb et al., 2022; Nguyen & Ene, 2024) for developing fast approximation algorithms for DSG has blossomed around applying convex-optimization algorithms for solving a tight linear programming (LP) relaxation of DSG originally proposed by Charikar (2000). The work of Danisch et al. (2017) considered a quadratic programming problem derived from the dual of the LP relaxation of Charikar (2000), and applied the Frank-Wolfe algorithm (Frank et al., 1956) for solving it. Recently, the authors of Harb et al. (2022) developed faster algorithms for DSG and dense decompositions on graphs by solving the dual LP relaxation via accelerated proximal gradient (Beck & Teboulle, 2009). This was later improved upon by Nguyen & Ene (2024) using tools from area convexity Sherman (2017) and accelerated random coordinate descent Ene & Nguyen (2015) to develop faster approximation algorithms for DSG and dense decompositions on graphs, respectively.
@@
An important variant of DSG is finding the $k$-core decomposition of a graph Seidman (1983). A $k$-core is the maximal subgraph (by size) where every vertex has an induced degree of at least $k$. The largest value of $k$ for which such a subgraph exists is known as the degeneracy, and the corresponding subgraph is known as the max-core. Extracting the max-core is tantamount to finding the vertex subset which maximizes the minimum induced degree. A slight modification of the greedy peeling algorithm for DSG exactly solves this problem.
@@
Recently, it has been shown (Veldt et al., 2021) that the solution of DSG and the $k$-core both correspond to special cases of a more general framework known as the Generalized Mean Densest Subgraph (GMDSG) problem. This framework is parameterized by a single parameter $p$, based on computing generalized means of degree sequences of a subgraph, with DSG and $k$-core corresponding to the choices of $p=1$ and $p=-\infty$, respectively. Hence, by varying $p$, one can extract a family of dense subgraphs which obey different notions of density. For $p\geq 1$, GMDSG can be solved optimally in polynomial-time via maximum flows, or approximated via a generalized greedy peeling algorithm. Additionally, it was shown by Chandrasekaran et al. (2024) that for $p<1$, GMDSG is NP–hard for $p$ in the range $p\in(-1/8,0)\cup(0,1/4)$, and that the solutions of DSG and max-core provide a $1/2$-approximation for GMDSG when $p<1$.
@@
Fairness in DSD: Subgraphs extracted via DSG or its variants are not guaranteed to fulfill a pre-specified fairness criterion (regardless of the algorithmic approach employed), as the respective formulations do not explicitly take into account any protected attributes that the vertices may possess. This issue was recently brought up by Anagnostopoulos et al. (2020; 2024); Miyauchi et al. (2023), which considered the task of extracting dense subgraphs that satisfy a group-fairness criterion. In Anagnostopoulos et al. (2020; 2024), each vertex is a member of either a protected or an unprotected group, and the goal is to extract the densest subgraph with an equal number of vertices from both groups (perfect balance). The follow-up work of Miyauchi et al. (2023) proposed a pair of fairness promoting formulations which can handle multiple vertex groups. The first (Miyauchi et al., 2023, Problem 1) corresponds to an extension of Anagnostopoulos et al. (2020; 2024) that aims to find the densest subgraph subject to the constraint that the number of vertices from each group in the solution does not exceed a fixed fraction of the solution size. The second formulation (Miyauchi et al., 2023, Problem 2) concerns finding the densest subgraph while ensuring that at least a pre-specified number of vertices from each group are present in the solution. A separate work (Oettershagen et al., 2024) considered a variation of the fair DSG problem where the edge relationships (as opposed to vertices) in a graph are modeled as belonging to different groups. The authors of Oettershagen et al. (2024) proposed formulations for extracting fair dense subgraphs containing exactly, at most, and at least a certain number of edges from each edge type.
@@ 3 Problem Statement
Recall the classic Densest Subgraph (DSG) problem (Goldberg, 1984). Given an undirected graph $\mathcal{G}$, DSG aims to extract a dense vertex subset $\mathcal{S}\subseteq\mathcal{V}$ that maximizes the average (induced) degree. The problem can be expressed as
Definition 1 (DSG).
${\mathbf{x}}^{*}=\operatorname*{argmax}_{{\mathbf{x}}\in\{0,1\}^{n}}~{}\rho({\mathbf{x}}),$ (1)
where $\rho({\mathbf{x}}):=({\mathbf{x}}^{\top}{\mathbf{A}}{\mathbf{x}})/({\mathbf{e}}^{\top}{\mathbf{x}})$ denotes the subgraph density of a candidate subgraph ${\mathbf{x}}\in\{0,1\}^{n}$ in terms of average induced-degree.
@@
Given a graph $\mathcal{G}$ with designated protected and unprotected subgroups of vertices, let ${\mathbf{e}}_{p}\in\{0,1\}^{n}$ denote the binary indicator vector of the protected subset $\mathcal{S}_{p}$. Define the regularizer $r_{1}({\mathbf{x}}):=({\mathbf{e}}_{p}^{\top}{\mathbf{x}})/{\mathbf{e}}^{\top}{\mathbf{x}}$ and consider the first Fairness-Aware Densest Subgraph (FADSG-I) problem.
Definition 2 (FADSG-I).
${\mathbf{x}}^{*}(\lambda)=\operatorname*{argmax}_{{\mathbf{x}}\in\{0,1\}^{n}}~{}\biggl{\{}g({\mathbf{x}},\lambda):=\rho({\mathbf{x}})+\lambda\cdot r_{1}({\mathbf{x}})\biggr{\}}.$ (2)
Here, $\lambda\geq 0$ is the regularization parameter. The objective function $g({\mathbf{x}},\lambda)$ consists of two terms - the first term $\rho({\mathbf{x}})$ corresponds to subgraph density, whereas for $\lambda>0$, $r_{1}({\mathbf{x}})$ promotes the inclusion of vertices from the protected subset $\mathcal{S}_{p}$. This can be seen by equivalently expressing $r_{1}(\cdot)$ as the set function
@@
It is evident that $0\leq r_{1}(\mathcal{S})\leq 1$, with the upper bound being attained for any subset $\mathcal{S}\subseteq\mathcal{S}_{p}$. Meanwhile, $r_{1}(\mathcal{S})=0$ iff $\mathcal{S}$ does not contain any protected vertices. Thus, larger values of $r_{1}(\cdot)$ correspond to vertex subsets containing a greater fraction of protected vertices. Note that $r_{1}(\mathcal{S})=1/2$ corresponds to the case of perfect balance (Anagnostopoulos et al., 2020; 2024), with equal number of protected and unprotected vertices in $\mathcal{S}$. Moreover, the choice of $r_{1}(\mathcal{S})=n_{p}/n$ implies that the number of vertices in the solution is proportional to the number of vertices in the population, a property known as demographic parity (Dwork et al., 2012). By varying $\lambda$, we allow $r_{1}(\cdot)$ to span a range of values, which affords flexibility in extracting dense subgraphs exhibiting varying fairness levels.
@@
When $\lambda=0$, FADSG-I reduces to DSG, the solution of which is not guaranteed to contain any protected vertices. As $\lambda$ is increased, greater emphasis is placed on maximizing $r_{1}({\mathbf{x}})$, enhancing the inclusion of protected vertices. For $\lambda\rightarrow\infty$, $r_{1}({\mathbf{x}}^{*}(\lambda))\rightarrow 1$, and the solution of problem (2) corresponds to the densest subset of $\mathcal{S}_{p}$. Clearly, by varying $\lambda$ in the interval $[0,\infty)$, the solution of (2) spans a spectrum which lies between these two extremes. By appropriately selecting $\lambda$, desired representation levels of the protected subset can be achieved. Hence, in contrast to Anagnostopoulos et al. (2020; 2024); Miyauchi et al. (2023), our group-fairness formulation (2) affords much greater flexibility in trading-off subgraph density in exchange for protected subset representation.
@@
It is worth pointing out that in certain scenarios, the solution of FADSG-I may not provide satisfactory levels of representation of the protected subset. Although the regularizer $r_{1}(\cdot)$ promotes the inclusion of protected vertices in the solution of (2), it does not explicitly take into account the fraction of such vertices that are represented. Even for (infinitely) large values of $\lambda$ in (2), the maximal level of representation attainable corresponds to the densest subset of $\mathcal{S}_{p}$. However, this does not imply that the fraction of the vertices of $\mathcal{S}_{p}$ that will be included in the solution of FADSG-I will be large (please see Table 7, Appendix H.3 for experimental results indicating this). These considerations motivate a more focused formulation for incorporating protected vertices in the solution, which we designate as the FADSG-II problem.
Definition 3 (FADSG-II).
${\mathbf{x}}^{*}(\lambda)\in\operatorname*{argmax}_{{\mathbf{x}}\in\{0,1\}^{n}}~{}\biggl{\{}h({\mathbf{x}},\lambda):=\rho({\mathbf{x}})-\lambda\cdot r_{2}({\mathbf{x}})\biggr{\}}.$ (4)
@@
In this case, $r_{2}({\mathbf{x}}):=(\lVert{\mathbf{x}}-{\mathbf{e}}_{p}\rVert_{2}^{2})/{\mathbf{e}}^{\top}{\mathbf{x}}$ is a regularizer and $\lambda\geq 0$ is a regularization parameter. The above problem differs from FADSG-I in the choice of the regularizer. Note that the numerator of $r_{2}(\cdot)$ can be viewed as the Hamming distance between the sets $\mathcal{S}$ and $\mathcal{S}_{p}$, since $\lVert{\mathbf{x}}-{\mathbf{e}}_{p}\rVert_{2}^{2}=\sum_{i\in\mathcal{V}}({\mathbf{x}}(i)-{\mathbf{e}}_{p}(i))^{2}$ corresponds to the number of disagreements between $\mathcal{S}$ (with indicator vector ${\mathbf{x}}$) and $\mathcal{S}_{p}$ (with indicator vector ${\mathbf{e}}_{p}$). Hence, for a given value of $\lambda>0$, the objective function of (4) aims to maximize subgraph density while penalizing the distance between the extracted and the protected subgraphs. As $\lambda\rightarrow\infty$, $r_{2}({\mathbf{x}}^{*}(\lambda))\rightarrow 0$, and the solution ${\mathbf{x}}^{*}(\lambda)$ coincides with the entire protected subset itself. Building on this observation, consider the case where the target fairness requirement stipulates that $50\%$ of the vertices in the protected subset be included in $\mathcal{S}$. We have the following result (see Appendix A.1 for proof).
@@
More generally, although problems (1) and (3) promote different notions of fairness, it is interesting to note that the two formulations are not completely unrelated. In particular, we have shown that the following relationship holds (see Appendix A.2 for the proof).
@@
For the regularizers $r_{1}(\mathcal{S})$ and $r_{2}(\mathcal{S})$, the following inequalities hold: $r_{1}(\mathcal{S})+r_{2}(\mathcal{S})\geq 1$ and $2r_{1}(\mathcal{S})+r_{2}(\mathcal{S})\leq 1+\frac{n_{p}}{2}$.
The implication of the above result is that given the value of one regularizer, we can attain bounds on the other (see Figure 9, Appendix A.2 for a visualization). Since providing explicit representation of the protected subset is absent in the fairness formulations of Anagnostopoulos et al. (2020; 2024); Miyauchi et al. (2023), FADSG-II goes beyond what is attainable using the prior art.
Tractability: A limitation of the formulations of Anagnostopoulos et al. (2020; 2024); Miyauchi et al. (2023) is that they are NP–hard in the worst-case. Since optimality cannot be guaranteed, this implies that the solution obtained by their approximation algorithms may not be satisfactory in terms of the loss in density incurred for a target fairness level. In contrast, it turns out that problems (2) and (4) can always be solved exactly in polynomial-time, irrespective of $\lambda$ (see Appendix B for details). This implies that by varying $\lambda$, we can trace the Pareto-optimal boundary between the two components comprising the objective function (i.e., density and fairness) of both problems (2) and (4). This enables us to reveal the underlying trade-off between density and fairness on real-world datasets. As mentioned before, the intrinsic hardness of the formulations in Anagnostopoulos et al. (2020; 2024); Miyauchi et al. (2023) does not allow for such a feature.